# The metric theory of small gaps for a sequence of real numbers

### v1

@@ 4.2. Proof of Theorem 6
where $\lfloor\widetilde{\delta}_{\min}^{\alpha}\rfloor(N)=\displaystyle\min_{1\leq m\neq n\leq N}\left\|\alpha\left\lfloor a_{m}-a_{n}\right\rfloor\right\|$, $\widetilde{B}_{N}=\displaystyle\max_{1\leq m\neq n\leq N}|\left\lfloor a_{m}-a_{n}\right\rfloor|$ and $H_{N}=\#\lfloor(A_{N}-A_{N})^{+}\rfloor$ are defined as in Section 2.1. In particular, $\widetilde{B}_{N}$ is upper bounded by $\max(\lfloor a_{1}\rfloor,\ldots,\lfloor a_{N}\rfloor)+1=\lfloor\max(a_{1},\ldots,a_{N})\rfloor+1$.
@@
(103) $\min_{1\leq m\neq n\leq N}\left\|\alpha\left(\lfloor a_{m}-a_{n}\right\rfloor+1)\right\|=\min_{1\leq m\neq n\leq N}\left\|\alpha\left(\lfloor a_{n}-a_{m}\right\rfloor+1)\right\|=$ $\min_{1\leq m\neq n\leq N}\left\|\alpha\left(-\lfloor a_{n}-a_{m}\right\rfloor-1)\right\|$
@@
(105) $\leq$ $\frac{4\log_{2}\widetilde{B}_{N}}{H_{N}\log(\frac{N}{9})\log_{2}(\frac{N}{9})},$
@@
Then, $\widetilde{B}_{N}\geq B_{N}$ always holds.
Furthermore, equality $\widetilde{B}_{N}=B_{N}$ holds if and only if the maximum value $\widetilde{B}_{N}$ is attained by a pair of indices $(m,n)$ for which the difference $|a_{m}-a_{n}|$ is an integer. In this case, $|\lfloor a_{m}-a_{n}\rfloor|=\lfloor|a_{m}-a_{n}|\rfloor$.
@@ 4.3. Proof of Theorem 9
Theorem 6 depends on the magnitude of $a_{N}$. To establish Theorem 9 (that is, to obtain a bound independent of $a_{N}$), we modify the coprime approach of (78) by removing only those subintervals centered at $a/Z_{n}$ such that $a$ and $Z_{n}$ share a small prime factor. This ensures the removed measure depends solely on $N$, not $a_{N}$, while sufficiently reducing problematic overlaps responsible for GCD sum growth.
@@ 4.3.1. Overlap estimates and auxiliary lemmas
To establish Theorem 9, it suffices to prove that almost every $\alpha$ belongs to infinitely many $S_{n}^{*}$. We now state the necessary lemmas, providing proofs for those without a cited reference.
@@
See [chandra2012borel, Theorem 1.4.3d] for a proof. ∎
Lemma 52 (Cassels’ zero-one law).
@@
Proof of Lemma 50. The following lemma adapts the overlap estimate of Pollington and Vaughan [pollingtonvaughan1990]. For notational convenience, we define throughout this proof
@@
Following Pollington and Vaughan’s method, we bound the contribution from distinct fractions $a/Z_{m}\neq b/Z_{n}$, noting that we only sifted primes below $4^{k}$. Let $u_{p}=\nu_{p}(Z_{m})$ and $v_{p}=\nu_{p}(Z_{n})$. Explicitly, we write
@@
where the sum over $j$ is estimated via Brun’s sieve, mirroring the technique employed by Pollington and Vaughan as in [pollingtonvaughan1990, p. 195–196].
@@
for some $j$ in $0\leq j\leq\gcd(Z_{m},Z_{n})$. We claim that for all $Z_{m}$ and $Z_{n}$ for which there exists a $p\leq 4^{k}$ such that the $p$-adic valuation of $Z_{m}$ is different from that of $Z_{n}$, there is no admissible value of $j$, and the double sum above is empty. In other words, a non-vanishing contribution requires that $\nu_{p}(Z_{m})=\nu_{p}(Z_{n})$ for all $p\leq 4^{k}$. However, if $\nu_{p}(Z_{m})=\nu_{p}(Z_{n})$ for all $p>4^{k}$ as well, then $Z_{m}=Z_{n}$, which is a contradiction. Therefore, a non-vanishing contribution implies that there exists a prime $p>4^{k}$ such that $\nu_{p}(Z_{m})\neq\nu_{p}(Z_{n})$.
We now prove the claim mentioned above by contrapositive. Suppose there exists an admissible value of $j$, meaning there exist $a$ and $b$ satisfying $0\leq a\leq Z_{m}$, $p\mid(a,Z_{m})\implies p>4^{k}$, and $0\leq b\leq Z_{n}$, $p\mid(b,Z_{n})\implies p>4^{k}$ such that $\frac{a}{Z_{m}}=\frac{b}{Z_{n}}$, that is, $aZ_{n}=bZ_{m}$. This shows that $\nu_{p}(Z_{m})=\nu_{p}(Z_{n})$ for all $p\leq 4^{k}$. If not, there exists $q\leq 4^{k}$ such that $\nu_{q}(Z_{m})\neq\nu_{q}(Z_{n})$. Without loss of generality, assume $\nu_{q}(Z_{n})>\nu_{q}(Z_{m})\geq 0$, that is, $\nu_{q}(Z_{n})-\nu_{q}(Z_{m})\geq 1$. Then $q$ divides $aq^{\nu_{q}(Z_{n})-\nu_{q}(Z_{m})}\displaystyle\prod_{p\neq q}p^{\nu_{p}(Z_{n})}=b\displaystyle\prod_{p\neq q}p^{\nu_{p}(Z_{m})}$, that is, $q$ divides $b$. Also, $\nu_{q}(Z_{n})\geq 1$ implies $q|Z_{n}$, so $q|(b,Z_{n})$ with $q\leq 4^{k}$, which is a contradiction. This contradiction shows that $\nu_{p}(Z_{m})=\nu_{p}(Z_{n})$ for all $p\leq 4^{k}$.
@@
Let $k$ be a fixed, “large”, even constant. We decompose all numbers $Z_{n}$ as follows.
@@
where $Z_{n}^{\text{smooth }}$ has only prime factors of size at most $4^{k}$, and $Z_{n}^{\text{rough }}$ only has prime factors of size larger than $4^{k}$, that is, $Z_{n}^{\text{smooth }}=\displaystyle\prod_{p\leq 4^{k}}p^{\nu_{p}(Z_{n})}$ and $Z_{n}^{\text{rough }}=\displaystyle\prod_{p>4^{k}}p^{\nu_{p}(Z_{n})}$. Write
@@
where $h$ is the uniquely defined index for which $Z_{n}^{\text{rough }}=b_{h}$, and where $p_{1}^{(k)}$ denotes the smallest prime exceeding $4^{k},p_{2}^{(k)}$ denotes the second-smallest prime exceeding $4^{k}$, and so on. The point in the construction of the numbers $\left(Y_{n}\right)_{2^{k/2}<n\leq 2^{k}}$ is that on the one hand in Lemma 50, we can replace $P(Z_{m},Z_{n})$ and $D(Z_{m},Z_{n})$ by $P\left(Y_{m},Y_{n}\right)$ and $D\left(Y_{m},Y_{n}\right)$ in the relevant situations, since the small prime factors of $Z_{m}$ and $Z_{n}$ are the same as those of $Y_{m}$ and $Y_{n}$, respectively. On the other hand, we have
@@
where we used that by construction the “small” prime factors of $Z_{n}$ and $Y_{n}$ coincide, and $Y_{n}$ has one additional prime factor which exceeds $4^{k}$. Consequently
@@
Thus, we can control the size of the Euler totient function of $Y_{n}$ (while we cannot control it for $Z_{n}$, which might have many large prime factors). This will allow us a straightforward application of Lemma 53. By Lemma 51, we have
@@
Note that whenever $Z_{m}^{\text{rough }}\neq Z_{n}^{\text{rough }}$, then we have $\gcd(Z_{m},Z_{n})\leq 4^{-k}\min(Z_{m},Z_{n})$ and so
$D(Z_{m},Z_{n})=\frac{\max\left(Z_{m}\psi(n),Z_{n}\psi(m)\right)}{\gcd(Z_{m},Z_{n})}\geq\frac{4^{k}\max(Z_{m},Z_{n})}{2^{k}\min(Z_{m},Z_{n})}\geq 2^{k},$
so that in this case
@@
We split the second sum in (125) according to whether $\left(Y_{m},Y_{n}\right)$ is in $\mathcal{E}_{4^{t+1}}$ or not. Bounding the Euler product trivially, we obtain
@@
(134) $\sum_{t=0}^{k-1}\sum_{\begin{subarray}{c}2^{k/2}<m,n\leq\leq 2^{k},\\ D\left(Y_{m},Y_{n}\right)\leq 4^{t+1},\\ (Y_{m},Y_{n})\in\mathcal{E}_{4}t+1\end{subarray}}\frac{\psi(m)\varphi\left(Y_{m}\right)}{Y_{m}}\frac{\psi(n)\varphi\left(Y_{n}\right)}{Y_{n}}\prod_{\begin{subarray}{c}p\mid\frac{Y_{m}Y_{n}}{\left(Y_{m},Y_{n}\right)^{2}}\\ p>4^{t+1}\end{subarray}}\left(1+\frac{1}{p}\right)$
(135) $\ll$ $\sum_{t=0}^{k-1}\frac{t}{4^{t}}$
(136) $\ll$ $1.$
@@
which in turn shows that
@@ 4.4. Proof of Theorem 11
To prove Theorem 11, we use methodology developed by Rudnick [Rudnick]. Let $A_{N}$ be defined as in Section 2.1. Denote the set of floored positive differences by $\mathcal{Z}_{N}$, that is, $\mathcal{Z}_{N}=\lfloor(A_{N}-A_{N})^{+}\rfloor=\{Z_{1},\dots,Z_{H_{N}}\}$. Therefore, $|\mathcal{Z}_{N}|=H_{N}$ and the hypothesis ensures that $\mathcal{Z}_{N}\subseteq\mathbb{N}$.
@@
which is $1$-periodic and localised on the scale $1/M$.
@@
$\int_{0}^{1}e(\alpha(k_{1}v_{1}-k_{2}v_{2}))d\alpha=\mathds{1}_{\{0\}}(k_{1}v_{1}-k_{2}v_{2}),$
where $\mathds{1}_{E}(*)$ is the indicator function of $E$, that is, $\mathds{1}_{E}(x)$ is equal to 1 if $x\in E$ and 0 otherwise.
@@
The proof proceeds in three steps. First, we establish a relationship between $H_{N_{k+1}}$ and $H_{N_{k}}$ along some subsequence $N_{k}$. In the next step, we show that the statistic $D(N,M)$ concentrates around its mean for an appropriate choice of $M$. Then, we use this concentration to deduce the existence of small gaps.
@@
By the Borel–Cantelli lemma, for almost every $\alpha$ there exists $k_{0}(\alpha)$ such that
@@
We have shown that for almost every $\alpha$, there exists $k_{0}(\alpha)$ such that $D(N_{k},M_{k})(\alpha)>1$ for all $k\geq k_{0}(\alpha)$. It then follows from the definition of $D(N,M)(\alpha)$ (See (140)) that for each $k\geq k_{0}(\alpha)$, there exists some $z\in\mathcal{Z}_{N_{k}}$ such that $F_{M_{k}}(\alpha z)>0$, that is, there exists $j_{0}\in\mathbb{Z}$ such that $f(M_{k}(\alpha z+j_{0}))>0$.
@@
Let $N\geq\lfloor k_{0}(\alpha)^{2/\eta}\rfloor$. Then $N\in\big[\lfloor(k_{0}(\alpha)+m)^{2/\eta}\rfloor,\lfloor(k_{0}(\alpha)+m+1)^{2/\eta}\rfloor\big)$ for some integer $m\geq 0$, that is, $N\in[N_{k_{0}(\alpha)+m},N_{k_{0}(\alpha)+m+1})$. Since $H_{N}$ is increasing in $N$, we have $H_{N_{k_{0}(\alpha)+m}}\leq H_{N}\leq H_{N_{k_{0}(\alpha)+m+1}}$. Applying Lemma 55, we obtain for all $N\geq\lfloor k_{0}(\alpha)^{2/\eta}\rfloor$,

### v2

@@ 4.2. Proof of Theorem 6
where $\lfloor\widetilde{\delta}_{\min}^{\alpha}\rfloor(N)=\displaystyle\min_{1\leq m\neq n\leq N}\left\|\alpha\left\lfloor a_{m}-a_{n}\right\rfloor\right\|$, $\widetilde{B}_{N}=\displaystyle\max_{1\leq m\neq n\leq N}|\left\lfloor a_{m}-a_{n}\right\rfloor|$, and $H_{N}=\#\lfloor(A_{N}-A_{N})^{+}\rfloor$. Since $\widetilde{B}_{N}\leq\max(\lfloor a_{1}\rfloor,\ldots,\lfloor a_{N}\rfloor)+1=\lfloor\max(a_{1},\ldots,a_{N})\rfloor+1$, the above statement also holds when $\widetilde{B}_{N}$ is replaced by $\lfloor\max(a_{1},\ldots,a_{N})\rfloor+1$.
@@
(102) $\min_{1\leq m\neq n\leq N}\left\|\alpha\left(\lfloor a_{m}-a_{n}\right\rfloor+1)\right\|=\min_{1\leq m\neq n\leq N}\left\|\alpha\left(\lfloor a_{n}-a_{m}\right\rfloor+1)\right\|=$ $\min_{1\leq m\neq n\leq N}\left\|\alpha\left(-\lfloor a_{m}-a_{n}\right\rfloor)\right\|$
@@
(104) $\leq$ $\frac{4\log_{2}\widetilde{B}_{N}}{H_{N}\log(\frac{N}{9})\log_{2}(\frac{N}{9})}$
@@
Then $\widetilde{B}_{N}\geq B_{N}$ always holds.
Furthermore, the equality $\widetilde{B}_{N}=B_{N}$ is valid if and only if the maximum value $\widetilde{B}_{N}$ is achieved by a pair of indices $(m,n)$ for which the difference $|a_{m}-a_{n}|$ is an integer. In this case, $|\lfloor a_{m}-a_{n}\rfloor|=\lfloor|a_{m}-a_{n}|\rfloor$.
@@ 4.3. Proof of Theorem 9
Theorem 6 depends on the magnitude of $a_{N}$. To establish Theorem 9 (that is, to obtain a bound independent of $a_{N}$), we modify the coprime approach of (77) by removing only those subintervals centred at $a/Z_{n}$ such that $a$ and $Z_{n}$ share a small prime factor. This ensures that the removed measure depends solely on $N$, not $a_{N}$, while sufficiently reducing the problematic overlaps responsible for the growth of the GCD sum.
@@ 4.3.1. Overlap estimates and auxiliary lemmas
To establish Theorem 9, it suffices to prove that almost every $\alpha$ belongs to infinitely many $S_{n}^{*}$. We now state the necessary lemmas, providing proofs for those for which no reference is given.
@@
See [chandra2012borel, Theorem 1.4.3d]. ∎
Lemma 51 (Cassels’ zero-one law [Cassels_1950]).
@@
Proof of Lemma 49. The following lemma adapts the overlap estimate of Pollington and Vaughan [pollingtonvaughan1990]. For notational convenience, we use the following notation throughout this proof
@@
Following Pollington and Vaughan’s method, we bound the contribution from distinct fractions $a/Z_{m}\neq b/Z_{n}$, noting that we only sifted primes below $4^{k}$. Let $u_{p}=\nu_{p}(Z_{m})$ and $v_{p}=\nu_{p}(Z_{n})$. Therefore,
@@
where the sum over $j$ is estimated by Brun’s sieve, mirroring the technique employed by Pollington and Vaughan as in [pollingtonvaughan1990, p. 195–196].
@@
for some $j$ in $0\leq j\leq\gcd(Z_{m},Z_{n})$. We claim that for all $Z_{m}$ and $Z_{n}$ for which there exists $p\leq 4^{k}$ such that the $p$-adic valuation of $Z_{m}$ is different from that of $Z_{n}$, there is no admissible value of $j$, and the double sum above is empty. In other words, a non-vanishing contribution requires that $\nu_{p}(Z_{m})=\nu_{p}(Z_{n})$ for all $p\leq 4^{k}$. However, if $\nu_{p}(Z_{m})=\nu_{p}(Z_{n})$ for all $p>4^{k}$ as well, then $Z_{m}=Z_{n}$, which is a contradiction. Therefore, a non-vanishing contribution implies that there exists a prime $p>4^{k}$ such that $\nu_{p}(Z_{m})\neq\nu_{p}(Z_{n})$.
We now prove the claim mentioned above by contrapositive. Suppose that there exists an admissible value of $j$, which means that there exist $a$ and $b$ satisfying $0\leq a\leq Z_{m}$, $p\mid(a,Z_{m})\implies p>4^{k}$, and $0\leq b\leq Z_{n}$, $p\mid(b,Z_{n})\implies p>4^{k}$ such that $\frac{a}{Z_{m}}=\frac{b}{Z_{n}}$, that is, $aZ_{n}=bZ_{m}$. This shows that $\nu_{p}(Z_{m})=\nu_{p}(Z_{n})$ for all $p\leq 4^{k}$. If not, there exists $q\leq 4^{k}$ such that $\nu_{q}(Z_{m})\neq\nu_{q}(Z_{n})$. Without loss of generality, assume $\nu_{q}(Z_{n})>\nu_{q}(Z_{m})\geq 0$, that is, $\nu_{q}(Z_{n})-\nu_{q}(Z_{m})\geq 1$. Then $q$ divides $aq^{\nu_{q}(Z_{n})-\nu_{q}(Z_{m})}\displaystyle\prod_{p\neq q}p^{\nu_{p}(Z_{n})}=b\displaystyle\prod_{p\neq q}p^{\nu_{p}(Z_{m})}$, that is, $q$ divides $b$. In addition, $\nu_{q}(Z_{n})\geq 1$ implies $q|Z_{n}$, so $q|(b,Z_{n})$ with $q\leq 4^{k}$, which is a contradiction. This contradiction shows that $\nu_{p}(Z_{m})=\nu_{p}(Z_{n})$ for all $p\leq 4^{k}$.
@@
Assume that $k$ is a fixed, even, and “large” number. We decompose all numbers $Z_{n}$ as follows.
@@
where $Z_{n}^{\text{smooth }}$ has only prime factors of size at most $4^{k}$, and $Z_{n}^{\text{rough }}$ has only prime factors of size larger than $4^{k}$, that is, $Z_{n}^{\text{smooth }}=\displaystyle\prod_{p\leq 4^{k}}p^{\nu_{p}(Z_{n})}$ and $Z_{n}^{\text{rough }}=\displaystyle\prod_{p>4^{k}}p^{\nu_{p}(Z_{n})}$. Write
@@
where $h$ is the uniquely defined index for which $Z_{n}^{\text{rough }}=b_{h}$, and where $p_{1}^{(k)}$ denotes the smallest prime exceeding $4^{k},p_{2}^{(k)}$ denotes the second-smallest prime exceeding $4^{k}$, and so on. The point in the construction of the numbers $\left(Y_{n}\right)_{2^{k/2}<n\leq 2^{k}}$ is that on the one hand, in Lemma 49, we can replace $P(Z_{m},Z_{n})$ and $D(Z_{m},Z_{n})$ by $P\left(Y_{m},Y_{n}\right)$ and $D\left(Y_{m},Y_{n}\right)$ in the relevant situations, since the small prime factors of $Z_{m}$ and $Z_{n}$ are the same as those of $Y_{m}$ and $Y_{n}$, respectively. On the other hand, we have
@@
where we used that by construction the “small” prime factors of $Z_{n}$ and $Y_{n}$ coincide, and $Y_{n}$ has an additional prime factor that exceeds $4^{k}$. Consequently,
@@
Thus, we can control the size of the Euler totient function of $Y_{n}$ (while we cannot control it for $Z_{n}$, which might have many large prime factors). This will allow us to apply Lemma 52. By Lemma 50, we have
@@
Note that whenever $Z_{m}^{\text{rough }}\neq Z_{n}^{\text{rough }}$, we have $\gcd(Z_{m},Z_{n})\leq 4^{-k}\min(Z_{m},Z_{n})$, and so
$D(Z_{m},Z_{n})=\frac{\max\left(Z_{m}\psi(n),Z_{n}\psi(m)\right)}{\gcd(Z_{m},Z_{n})}\geq\frac{4^{k}\max(Z_{m},Z_{n})}{2^{k}\min(Z_{m},Z_{n})}\geq 2^{k}.$
Therefore, in this case
@@
We divide the second sum in the last line of (124) depending on whether $\left(Y_{m},Y_{n}\right)$ is in $\mathcal{E}_{4^{t+1}}$ or not. Bounding the Euler product trivially, we obtain
@@
(133) $\sum_{t=0}^{k-1}\sum_{\begin{subarray}{c}2^{k/2}<m,n\leq\leq 2^{k},\\ D\left(Y_{m},Y_{n}\right)\leq 4^{t+1},\\ (Y_{m},Y_{n})\in\mathcal{E}_{4}t+1\end{subarray}}\frac{\psi(m)\varphi\left(Y_{m}\right)}{Y_{m}}\frac{\psi(n)\varphi\left(Y_{n}\right)}{Y_{n}}\prod_{\begin{subarray}{c}p\mid\frac{Y_{m}Y_{n}}{\left(Y_{m},Y_{n}\right)^{2}}\\ p>4^{t+1}\end{subarray}}\left(1+\frac{1}{p}\right)\ll\sum_{t=0}^{k-1}\frac{t}{4^{t}}\ll 1.$
@@
which implies that
@@ 4.4. Proof of Theorem 11
To prove Theorem 11, we use the methodology developed by Rudnick [Rudnick]. Let $A_{N}$ be defined as in Section 2.1. Denote the set of floored positive differences by $\mathcal{Z}_{N}$, that is, $\mathcal{Z}_{N}=\lfloor(A_{N}-A_{N})^{+}\rfloor=\{Z_{1},\dots,Z_{H_{N}}\}$. Therefore, $|\mathcal{Z}_{N}|=H_{N}$ and the hypothesis ensure that $\mathcal{Z}_{N}\subseteq\mathbb{N}$.
@@
which is $1$-periodic and is localised on the scale $1/M$.
@@
$\int_{0}^{1}e(\alpha(k_{1}v_{1}-k_{2}v_{2}))d\alpha=\mathds{1}_{\{0\}}(k_{1}v_{1}-k_{2}v_{2}).$
@@
The proof is carried out in three steps. First, we establish a relationship between $H_{N_{k+1}}$ and $H_{N_{k}}$ along some subsequence $N_{k}$. In the next step, we show that the statistic $D(N,M)$ concentrates around its mean for an appropriate choice of $M$. Then, we use this concentration to deduce the existence of small gaps.
@@
By the Borel–Cantelli lemma, for almost all $\alpha$, there exists $k_{0}(\alpha)$ such that
@@
We have shown that for almost all $\alpha$, there exists $k_{0}(\alpha)$ such that $D(N_{k},M_{k})(\alpha)>1$ for all $k\geq k_{0}(\alpha)$. It then follows from the definition of $D(N,M)(\alpha)$ (see (137)) that for each $k\geq k_{0}(\alpha)$, there exists some $z\in\mathcal{Z}_{N_{k}}$ such that $F_{M_{k}}(\alpha z)>0$, that is, there exists $j_{0}\in\mathbb{Z}$ such that $f(M_{k}(\alpha z+j_{0}))>0$.
@@
Let $N\geq\lfloor k_{0}(\alpha)^{2/\eta}\rfloor$. Then $N\in\big[\lfloor(k_{0}(\alpha)+m)^{2/\eta}\rfloor,\lfloor(k_{0}(\alpha)+m+1)^{2/\eta}\rfloor\big)$ for some integer $m\geq 0$, that is, $N\in[N_{k_{0}(\alpha)+m},N_{k_{0}(\alpha)+m+1})$. Since $H_{N}$ increases in $N$, we have $H_{N_{k_{0}(\alpha)+m}}\leq H_{N}\leq H_{N_{k_{0}(\alpha)+m+1}}$. Applying Lemma 54, we obtain for all $N\geq\lfloor k_{0}(\alpha)^{2/\eta}\rfloor$,